# Frequency dependence of nonsequential double ionization of atoms in strong laser fields

### v1

@@ V Conclusion
Despite the continuous development of numerical methods and the growth of computer resources, the solving of full–dimensionall quantum mechanical problem of atoms interacting with strong short laser pulses is still a big challenge. The scanning of a wide range of frequency (wavelength) and laser pulse intensity when calculating ionization yields requires the use of a significant amount of calculation time resources. Going further, the generation of other experimental observables, e.g. momentum distributions, sets high requirements on the operating memory and disk resources, transferring the center of gravity towards the big data analysis. While the solutions in this area exist, they will not always be a method of first choice. When the compromise threshold between the experiment and the theoretical description is determined at the level of qualitative explanations of the observed phenomenon, simplified models become fully sufficient. Indeed, this is the case of the problem investigated by us. With use of a quantum mechanical (1+1)-dimensional model we analyzed the frequency dependence of strong-field double ionization with special regard to experimentally available data, i.e. intensity dependent single and double ionization yields, ratios of double to single ionization, electrons momenta distributions and longitudinal ion momentum distributions.
@@
The obtained dependence of the single and double ionization yields on the laser frequency is in agreement with available experimental data and semi-classical calculations. In our approach we used He-like simplified model, however, accessible data concern Ar, He, Mg, Ne and Xe atoms. The observed agreement, once again, proves that our model catches the essence of the analyzed phenomena, despite of reduced dimensionality. Usually, decrease in the ratios of double to single ions with drecreasing $\omega$ is attributed to interplay between wave-packet diffusion, Coulomb focusing and defocusing effects. However, the later two cannot be accounted for by a one-dimensional model and the wave-packet diffusion is also strongly suppressed in reduced-dimensions. The observed change in the ratio of double to single ions is attributed the smaller role played by RESI at lower frequencies. Next, the influence of pulse duration on the obtained yields was analyzed showing that knee regime is shifted towards lower laser intensities in longer pulses. The obtained results were corroborated by considering rate equations and exploiting the fact that the ratio of double to single ionization is almost constant in the knee regime.
Finally, the study is supplemented with two-electron momentum distributions and longitudinal ion momentum distributions. The transition form nonsequential to sequential regimes is clearly visible for all studied laser frequencies. Increase of the momentum space in the momentum distributions with decreasing $\omega$ (increasing $F_{0}$) is also observed. The smaller role played by RESI at lower frequencies is confirmed by the observed changes in the longitudinal ion momentum distributions with decreasing frequency.
@@ Acknowledgements.
J.P-B.: A tribute to the late Bruno Eckhardt with whom I had the honor of working. The above work was initiated and stimulated by him.

### v2

@@ V Conclusion
Despite the continuous development of numerical methods and the growth of computer resources, solving the full–dimensional quantum mechanical problem of atoms interacting with strong short laser pulses is still a big challenge. Scanning a wide range of laser frequencies (wavelengths) and intensities when calculating ionization yields requires a significant amount of computational resources. Going further, the generation of other experimental observables, e.g. momentum distributions, sets high requirements on the computer memory and disk resources, shifting the focus to the big data analysis. While solutions in this area exist, they will not always be the first method of choice. When the line of compromise between the experiment and the theoretical description is determined at the level of qualitative explanations of the observed phenomenon, simplified models are fully sufficient.
@@
Indeed, this is the case of the problem investigated by us. Using a quantum mechanical (1+1)-dimensional model, we analyzed the frequency dependence of strong-field double ionization with special regard to experimentally available data, i.e. intensity-dependent single and double ionization yields, ratios of double to single ionization, electron momentum distributions and longitudinal ion momentum distributions. The observed dependence of the single and double ionization yields on the laser frequency is in agreement with available experimental data and semi-classical calculations. In our approach, we used a simplified He-like model, however, accessible data concerns Ar, He, Mg, Ne and Xe atoms. The observed agreement, once again, proves that our model catches the essence of the analyzed phenomena, despite the reduced dimensionality. Usually, the decrease in the ratios of double to single ions with decreasing $\omega$ is attributed to the interplay of wave-packet diffusion, Coulomb focusing and defocusing effects. However, the latter two cannot be accounted for by a reduced-dimensionality model, and wave-packet diffusion is also strongly suppressed there. The observed change in the ratio of double to single ionization is attributed to the smaller role played by RESI at lower frequencies. Next, the influence of the pulse duration on the obtained yields was analyzed, showing that the knee regime is shifted towards lower laser intensities in longer pulses. The obtained results were corroborated by considering rate equations and exploiting the fact that the ratio of double to single ionization is almost constant in the knee regime. Finally, the study was supplemented with two-electron momentum distributions and longitudinal ion momentum distributions. The transition from the nonsequential to the sequential regime is clearly visible for all studied laser frequencies. The increase of the momentum space in the momentum distributions with decreasing $\omega$ and increasing $F_{0}$ is also observed. The smaller role played by RESI at lower frequencies is confirmed by the observed changes in the longitudinal ion momentum distributions with decreasing frequency.
@@ Acknowledgements.
A tribute to the late Bruno Eckhardt with whom we had the honor of working. The above work was initiated and stimulated by him.